OpenStudy (anonymous):

Solve the equation for y: 4x + 5y + 15 = 10.

OpenStudy (anonymous):

4x + 5y + 15 - 10 = 0 4x + 5y + 5 = 0 5y = -4x - 5 y = (-4x - 5)/5

OpenStudy (ajprincess):

4x + 5y + 15 = 10. 1.Subtract 15 both sides. 4x+5y+15-15=10-15 4x+5y=-5 2.Subtract 4x both sides 4x+5y-4x=-5-4x 5y=-5-4x 3.Divide by 5 both sides 5y/5=(-5-4x)/5 y=(-5-4x)/5 y=-(5+4x)/5 Is that clear?

OpenStudy (anonymous):

Yes . thank you!

OpenStudy (ajprincess):

yw.:)